# Symmetry-Constrained Forecasting of Periodically Correlated Energy Processes

### v1

@@ Abstract
Time series in energy systems, such as solar irradiance, wind speed, or electrical load, are characterized by strong diurnal and seasonal periodicities. Accurate forecasting requires accounting for time varying statistical properties that stationary or classical persistence models cannot capture. A family of analytical forecasting operators for cyclostationary processes is introduced, extending persistence through a closed form coefficient $\tilde{\lambda}(t,\tau)=\tfrac{1}{2}\bigl(1+\rho(t,\tau)\bigr)$, where $\rho(t,\tau)$ denotes the local correlation between the current observation and its phase aligned counterpart. This formulation preserves periodic variance and covariance, achieving a symmetry induced reduction of effective degrees of freedom. The resulting operator defines a training free analytical limit of persistence under periodic non stationarity. Validation on synthetic cyclostationary signals and empirical renewable energy datasets demonstrates consistent accuracy gains over classical persistence, particularly at multi hour horizons. By embedding temporal symmetry into the prediction process, the framework provides a physically interpretable, reproducible, and computationally minimal baseline for forecasting periodic processes across energy and complex systems.
@@
journal: Engineering Applications of Artificial Intelligence
@@ 1 Introduction
Time series in energy systems, including solar irradiance, wind speed, and electrical demand, exhibit strong daily and seasonal periodicities arising from astronomical and human cycles [18]. Accurate forecasting of these signals is essential for grid stability, renewable integration, and energy market operation [20, 8]. However, the underlying processes are intrinsically non-stationary: both mean and variance evolve with time in a nearly periodic manner. Most forecasting models (from classical statistical approaches to deep learning architectures) implicitly assume stationarity, capturing short-term dependencies while neglecting the periodic modulation of statistical moments [33]. This structural mismatch limits both interpretability and generalization in operational contexts [22].
@@
Recent computational developments (notably deep learning and hybrid machine learning methods) have shown strong potential for modeling complex temporal structures in renewable-energy time series [11, 15]. Yet these models remain constrained by extensive training-data requirements and high computational costs, which limits real-time applicability [24, 2]. Analytical formulations based on cyclostationary processes provide an alternative by explicitly accounting for the periodic variation of statistical moments [4, 27]. Periodic autoregressive ($\mathtt{PAR}$) models and exponential-smoothing families such as Holt–Winters ($\mathtt{HW}$) or Error–Trend–Seasonal ($\mathtt{ETS}$) [31, 5], together with more advanced frameworks like Theta [30], Prophet [1], $\mathtt{N\text{-}BEATS}$, and $\mathtt{TBATS}$ [35, 12], can achieve high accuracy but at substantial computational complexity, limiting deployment in operational environments where efficiency and interpretability are essential.
@@
Table 1: Forecasting methods for periodic time series. $X_{t}$ is the measured value at time $t$; $n\Delta t$ is the forecast horizon; $T$ is the seasonal period; $p$ is the AR order; $k$ the number of Theta lines; $B$ the block count; $d$ the network depth in $\mathtt{N\text{-}BEATS}$; $m$ the number of seasonal components in $\mathtt{TBATS}$. For $\mathtt{HW}$ and $\mathtt{ETS}$, $L(t)$, $b(t)$, and $S(t)$ are level, trend, and seasonal components. For Prophet, $g(t)$, $s(t)$, and $h(t)$ are the growth, seasonal, and holiday effects. For $\mathtt{SARIMA}$, $\phi(B)$ and $\Phi(B^{T})$ are AR polynomials, $(1-B)^{d}$ and $(1-B^{T})^{D}$ are differencing operators, $\theta(B)$ and $\Theta(B^{T})$ are $\mathtt{MA}$ polynomials, and $\epsilon(t)$ is white noise. For $\mathtt{N\text{-}BEATS}$, $\theta_{i}$ are trainable parameters and $b_{i}(t)$ basis functions; for $\mathtt{TBATS}$, $s(i,t)$ denotes seasonal components.
Method Forecast Equation for $\hat{X}(t+n\Delta t)$ $\mathcal{C}_{T}$ $\mathcal{C}_{F}$
Holt-Winters ($\mathtt{HW}$) $L(t)+b(t)n\Delta t+S(t-T+(n\Delta t\bmod T))$ $\mathcal{O}(n)$ $\mathcal{O}(h)$
@@
Error-Trend-Seasonal ($\mathtt{ETS}$) $L(t)+b(t)n\Delta t+S(t-T+(n\Delta t\bmod T))$ $\mathcal{O}(n)$ $\mathcal{O}(h)$
@@
Theta Method $\theta X(t)+(1-\theta)\bigl(L(t)+b(t)n\Delta t\bigr)$ $\mathcal{O}(nk)$ $\mathcal{O}(h)$
Periodic $\mathtt{AR}$ ($\mathtt{PAR}$) $\sum_{i=1}^{p}a_{i}\,X(t+n\Delta t-i)$ $\mathcal{O}(np^{2}T)$ $\mathcal{O}(p)$
Cyclostationary $\mathtt{AR}$ ($\mathtt{CAR}$) $\sum_{i=1}^{p}a_{i}\bigl((t+n\Delta t)\bmod T\bigr)\,X(t+n\Delta t-i)$ $\mathcal{O}(np^{2}T)$ $\mathcal{O}(pT)$
Prophet $g(t+n\Delta t)+s(t+n\Delta t)+h(t+n\Delta t)$ $\mathcal{O}(n\log n+nk)$ $\mathcal{O}(T)$
$\mathtt{SARIMA}$ $\phi(B)\,\Phi(B^{T})\,(1-B)^{d}(1-B^{T})^{D}X(t)+\theta(B)\,\Theta(B^{T})\,\epsilon(t)$ $\mathcal{O}(np^{3})$ $\mathcal{O}(p)$
$\mathtt{STL}$+ $\mathtt{ETS}$ $T(t)+S\bigl(t+(n\Delta t\bmod T)\bigr)+R(t)$ $\mathcal{O}(n\log n)$ $\mathcal{O}(T+n\Delta t)$
$\mathtt{TBATS}$ $l(t)+\sum_{i=1}^{m}s\bigl(i,t+n\Delta t\bigr)$ $\mathcal{O}(nm^{3})$ $\mathcal{O}(m^{2})$
$\mathtt{N\text{-}BEATS}$ $\sum_{i=1}^{B}\theta_{i}\,b_{i}\bigl(t+n\Delta t\bigr)$ $\mathcal{O}(nBd)$ $\mathcal{O}(Bd)$
@@
Table 1 summarizes principal forecasting models incorporating periodic or seasonal components, providing for each method the predictive formulation $\hat{X}(t+n\Delta t)$ and the computational complexity of training and inference. This overview establishes quantitative benchmarks for evaluating new forecasting operators under periodic non-stationarity. In an era dominated by increasingly complex data-driven architectures, the present work instead pursues analytical parsimony: seeking understanding over computation through an interpretable formulation that requires neither exogenous inputs nor substantial computational resources.
@@ 1.1 Motivation and Scope
According to [9], Persistence (the assumption that future values equal current observations) remains the operational baseline of renewable energy forecasting [33]; however, its performance degrades when periodic nonstationary effects modulate the variance and correlation structure, a behavior documented in cyclic environmental processes. Its simplicity, absence of calibration, and negligible computational cost have made it the universal benchmark for deterministic prediction [3, 24]. However, for renewable processes such as solar irradiance ($I$) or wind speed, whose mean, variance, and correlation vary systematically with daily and seasonal cycles, classical persistence rapidly loses validity beyond intra-hour horizons. Reviews across solar and wind forecasting consistently highlight both the methodological centrality and the structural limitations of persistence based approaches under periodic non-stationarity [19, 16].
Several persistence variants have been proposed to mitigate these effects. The Complete-History Persistence Ensemble ($\mathtt{CH\!-\!PeEn}$) aggregates past observations at identical daily phases [36]. Smart Persistence incorporates clear-sky modeling but remains sensitive to meteorological inputs and model mismatch [28, 26]. Clearness index formulations improve robustness and computational efficiency [23]. Numerous training free forecasting approaches have been proposed, including nearest-neighbor and functional data methods designed to improve persistence baselines. These methods primarily target predictive accuracy under specific operational contexts [21]. The $\mathtt{CLIPER}$ method establishes an $\mathtt{MSE}$ optimal balance between climatology and persistence, providing a stationary analytical baseline [37, 39]. Yet all these approaches rely on quasi-stationarity assumptions and therefore cannot preserve the cyclic evolution of second-order statistics that characterizes renewable signals. A complementary transformer-based semantic screening of recent forecasting literature (all-MiniLM-L6-v2 applied to Scopus, CrossRef, and OpenAlex) confirms this tendency: among 86 publications (2010–2024) addressing persistence based forecasting, only 37% analyzed persistence explicitly and 14% proposed methodological extensions. Publication counts exhibit exponential growth ($R^{2}=0.82$), with time series models representing 49% of works, numerical weather prediction approaches 30%, and satellite-based methods 19%. This indicates that persistence remains primarily treated as a benchmark rather than as a methodological subject.
@@ 1.2 Contributions and Structure of the Paper
Table 2: Summary of the forecasting operators introduced in this study, with their analytical regimes, computational complexities, and governing parameters. Each operator acts as a linear mapping characterized by training and forecasting complexities $\mathcal{C}_{T};\mathcal{C}_{F}]$ as described in App. A.
@@ 2.1 Classical Models
Definition 1 (Simple Persistence Operator $\emptyset;\mathcal{O}(1)]$).
@@
Definition 2 (Cyclic Persistence Operator $\emptyset;\mathcal{O}(1)]$).
@@
Definition 3 (Smart Persistence Operator $\mathcal{O}(\mathtt{ref});\mathcal{O}(1)]$).
@@
Definition 4 ($\mathtt{CLIPER}$ Operator under Stationary Assumptions $\mathcal{O}(n);\mathcal{O}(1)]$).
The $\mathtt{CLIPER}$ operator $\mathtt{P}_{\mathtt{CLIPER}}$ predicts a future value as a convex combination of the process mean $\mu$ and the most recent observation $I(t)$. Analytical formulations of this operator in stationary contexts are discussed in [37, 32]. It is defined as
@@
For $0\leq\rho(n\Delta t)\leq 1$, $\lambda^{\mathtt{CLIPER}}$ lies within $[0,1]$, yielding a valid convex combination between mean and current value. When $\rho(n\Delta t)<0$, the parameter may exceed unity, reflecting anti-correlated regimes where the mean contribution dominates. The complete derivation from $\mathtt{MSE}$ minimization is provided in App. B. Cyclostationary models provide a natural framework for processes exhibiting periodic statistical structure, allowing predictors to remain consistent with cyclic variations in mean and covariance.
@@ 2.2 Cyclostationary Extension of Forecasting Operators
Theorem 1 (Periodically Stationary or Cyclostationary Processes).
@@
In contrast to weakly stationary processes, cyclostationary signals preserve their statistical structure modulo the period $T$. Their moments evolve periodically, reflecting temporal symmetries common to many natural and engineered systems such as irradiance, wind speed, electrical load, or tidal dynamics. This property defines the statistical foundation required for analytical forecasting under periodic non-stationarity.
Statistical Stationary Process Cyclostationary Process
$E[X(t)]$ $\mu(t)=\mu,\ \forall t$ $\mu(t)=\mu(t+T),\ \forall t$
@@
The forecasting workflow using $\mathtt{P}_{\mathtt{BLEND}}$ is summarized in Figure 1, which distinguishes the stationary, periodically stationary, and simplified analytical forms. The mathematical properties detailed in Appendix D ensure convexity, boundedness, and stability, core guarantees for applying $\mathtt{BLEND}$ operators to cyclostationary energy processes where periodicity and noise coexist.

### v3

@@ Abstract
Time series in energy systems, such as solar irradiance, wind speed, or electrical load, are characterized by strong diurnal and seasonal periodicities. Accurate forecasting requires accounting for time varying statistical properties that stationary or classical persistence models cannot capture. A family of analytical forecasting operators for cyclostationary processes is introduced, extending persistence through a closed form coefficient $\tilde{\lambda}(t,\tau)=\tfrac{1}{2}\bigl(1+\rho(t,\tau)\bigr)$, where $\rho(t,\tau)$ denotes the local correlation between the current observation and its phase aligned time lag ($\tau$). This formulation preserves periodic variance and covariance, achieving a symmetry induced reduction of effective degrees of freedom. The resulting operator defines a training free analytical limit of persistence under periodic non stationarity. Validation on synthetic cyclostationary signals and empirical renewable energy datasets demonstrates consistent accuracy gains over classical persistence, particularly at multi hour horizons. By embedding temporal symmetry into the prediction process, the framework provides a physically interpretable, reproducible, and computationally minimal baseline for forecasting periodic processes across energy and complex systems.
@@
journal: Applied Mathematical Modelling
@@ 1 Introduction
Time series in energy systems, including solar irradiance, wind speed, and electrical demand, exhibit strong daily and seasonal periodicities arising from astronomical and human life cycles [21]. Accurate forecasting of these signals is essential for grid stability, renewable integration, and energy market operation [24, 11]. However, the underlying processes are intrinsically non-stationary: both mean and variance evolve with time in a nearly periodic manner. Most forecasting models (from classical statistical approaches to deep learning architectures) implicitly assume stationarity, capturing short-term dependencies while neglecting the periodic modulation of statistical moments [39]. This structural mismatch limits both interpretability and generalization in operational contexts [26].
@@
Recent computational developments (notably deep learning and hybrid machine learning methods) have shown strong potential for modeling complex temporal structures in renewable-energy time series [15, 18]. Yet these models remain constrained by extensive training-data requirements and high computational costs, which limits real-time applicability [28, 3]. Analytical formulations based on cyclostationary processes provide an alternative by explicitly accounting for the periodic variation of statistical moments [13, 31]. Periodic autoregressive ($\mathtt{PAR}$) models and exponential-smoothing families such as Holt–Winters ($\mathtt{HW}$) or Error–Trend–Seasonal ($\mathtt{ETS}$) [35], together with more advanced frameworks like Theta [2], Prophet [1], $\mathtt{N\text{-}BEATS}$, and $\mathtt{TBATS}$ [9], can achieve high accuracy but at substantial computational complexity, limiting deployment in operational environments where efficiency and interpretability are essential. Table 1 summarizes representative forecasting models with periodic or seasonal structure. For each method, a simplified predictive expression for $\hat{X}(t+n\Delta t)$ together with approximate training and forecasting complexities is reported for comparative purposes.
@@
Table 1: Simplified forecasting equations for periodic time series. $X_{t}$ is the measured value at time $t$; $n\Delta t$ is the forecast horizon ($n\geq 1$); $T$ is the seasonal period; $p$ is the $\mathtt{AR}$ order; $k$ the number of Theta lines; $N_{B}$ the block count; $d$ the network depth in $\mathtt{N\text{-}BEATS}$ and $\theta_{i}$ denote trainable parameters while $b_{i}(t)$ learned basis functions defined by the network architecture; for $\mathtt{TBATS}$, $s(i,t)$ denotes model-specific seasonal components and $m$ the number of seasonal components. For $\mathtt{HW}$, $\mathtt{ETS}$ and $\mathtt{TBATS}$, $L(t)$, $b(t)$, and $S(t)$ are level, trend, and seasonal components while $\Phi_{n\Delta t}$ is the cumulative damping factor of the trend over the horizon $n\Delta t$. For $\mathtt{STL}$, $Tr(t)$ denotes the trend component obtained via local regression smoothing and the expression is a simplified reconstruction. For Prophet, $g(t)$, $s(t)$, and $h(t)$ are the growth, seasonal, and holiday effects. For $\mathtt{SARIMA}$ (and $\mathtt{PAR}$), $\phi(B)$ and $\Phi(B^{T})$ are $\mathtt{AR}$ polynomials, $(1-B)^{d}$ and $(1-B^{T})^{D}$ are differencing operators, $\theta(B)$ and $\Theta(B^{T})$ are $\mathtt{MA}$ polynomials, and $\epsilon(t)$ is white noise. Complexities $\mathcal{C}_{T}$ and $\mathcal{C}_{F}$ denote approximate orders of magnitude for training and forecasting costs, reported for comparative purposes only (see A).
Method Approximate Equation for $\hat{X}(t+n\Delta t)$ $\mathcal{C}_{T}$ $\mathcal{C}_{F}$
Holt-Winters ($\mathtt{HW}$) [6] $L(t)+n\Delta t.b(t)+S(t-T+n\Delta t)$ $\mathcal{O}(n)$ $\mathcal{O}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n})$
@@
Error-Trend-Seasonal ($\mathtt{ETS}$) [23] $L(t)+\Phi_{n\Delta t}.b(t)+S(t-T+n\Delta t)$ $\mathcal{O}(n)$ $\mathcal{O}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n})$
@@
Theta Method [2] $\theta X(t)+(1-\theta)\bigl(L(t)+b(t)n\Delta t\bigr)$ $\mathcal{O}(nk)$ $\mathcal{O}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n})$
Periodic $\mathtt{AR}$ ($\mathtt{PAR}$) [36] $\sum_{i=1}^{p}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{i}(t\bmod T)}\,X(t+n\Delta t-i)$ $\mathcal{O}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}np^{2}})$ $\mathcal{O}(p)$
Prophet [1] $g(t+n\Delta t)+s(t+n\Delta t)+h(t+n\Delta t)$ $\mathcal{O}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}nk})$ $\mathcal{O}(n)$
$\mathtt{SARIMA}$ [4] $\phi(B)\,\Phi(B^{T})\,(1-B)^{d}(1-B^{T})^{D}X(t)=\theta(B)\,\Theta(B^{T})\,\epsilon(t)$ $\mathcal{O}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}np^{2}})$ $\mathcal{O}(p)$
$\mathtt{STL}$ [7] ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Tr(t+n\Delta t)}+S\bigl(t+(n\Delta t\bmod T)\bigr)$ - -
$\mathtt{TBATS}$ [9] $L(t)+\sum_{i=1}^{m}s\bigl(i,t+n\Delta t\bigr)$ - -
$\mathtt{N\text{-}BEATS}$ [32] $\sum_{i=1}^{N_{B}}\theta_{i}\,b_{i}\bigl(t+n\Delta t\bigr)$ - -
For $\mathtt{N\text{-}BEATS}$, $\mathtt{STL}$ and $\mathtt{TBATS}$, computational complexities are not reported, as they depend on model architecture, parameterization, implementation choices and no unique asymptotic expression can be defined. These models are therefore excluded from the comparative complexity analysis.
@@
This overview establishes quantitative benchmarks for evaluating new forecasting operators under periodic non stationarity. In an era dominated by increasingly complex data-driven architectures, the present work instead pursues analytical parsimony: seeking understanding over computation through an interpretable formulation that requires neither exogenous inputs nor substantial computational resources.
@@ 1.1 Motivation and Scope
According to [12], Persistence (the assumption that future values equal current observations) remains the operational baseline of renewable energy forecasting [39]; however, its performance degrades when periodic nonstationary effects modulate the variance and correlation structure, a behavior documented in cyclic environmental processes. Its simplicity, absence of calibration, and negligible computational cost have made it the universal benchmark for deterministic prediction [28, 5]. However, for renewable processes such as solar irradiance ($I$) or wind speed, whose mean, variance, and correlation vary systematically with daily and seasonal cycles, classical persistence rapidly loses validity beyond intra-hour horizons. Reviews across solar and wind forecasting consistently highlight both the methodological centrality and the structural limitations of persistence based approaches under periodic non stationarity [22, 19].
Several persistence variants have been proposed to mitigate these effects. The Complete-History Persistence Ensemble ($\mathtt{CH\!-\!PeEn}$) aggregates past observations at identical daily phases [41]. Smart Persistence incorporates clear-sky modeling but remains sensitive to meteorological inputs and model mismatch [33, 30]. Clearness index formulations improve robustness and computational efficiency [27]. Numerous training free forecasting approaches have been proposed, including nearest-neighbor and functional data methods designed to improve persistence baselines. These methods primarily target predictive accuracy under specific operational contexts [25]. The $\mathtt{CLIPER}$ method establishes an $\mathtt{MSE}$ optimal balance between climatology and persistence, providing a stationary analytical baseline [41]. Yet all these approaches rely on quasi-stationarity assumptions and therefore cannot preserve the cyclic evolution of second-order statistics that characterizes renewable signals. A complementary transformer-based semantic screening of recent forecasting literature (all-MiniLM-L6-v2 applied to Scopus, CrossRef, and OpenAlex) confirms this tendency: among 86 publications (2010–2024) addressing persistence based forecasting, only 37% analyzed persistence explicitly and 14% proposed methodological extensions. Publication counts exhibit exponential growth ($R^{2}=0.82$), with time series models representing 49% of works, numerical weather prediction approaches 30%, and satellite-based methods 19%. This indicates that persistence remains primarily treated as a benchmark rather than as a methodological subject.
@@ 1.2 Contributions and Structure of the Paper
Table 2: Summary of the forecasting operators introduced in this study, with their analytical regimes, computational complexities, and governing parameters. Each operator acts as a linear mapping characterized by training and forecasting complexities $\mathcal{C}_{T};\mathcal{C}_{F}]$ as described in  A.
@@ 2.1 Classical Models
Definition 1 (Simple Persistence Operator [12] $\emptyset;\mathcal{O}(1)]$).
@@
Definition 2 (Cyclic Persistence Operator [31] $\emptyset;\mathcal{O}(1)]$).
@@
Definition 3 (Smart Persistence Operator [28] $\mathcal{O}(\mathtt{ref});\mathcal{O}(1)]$).
@@
Definition 4 ($\mathtt{CLIPER}$ Operator under Stationary Assumptions [41] $\mathcal{O}(n);\mathcal{O}(1)]$).
The $\mathtt{CLIPER}$ operator $\mathtt{P}_{\mathtt{CLIPER}}$ predicts a future value as a convex combination of the process mean $\mu$ and the most recent observation $I(t)$. Analytical formulations of this operator in stationary contexts are discussed in [38]. It is defined as
@@
For $0\leq\rho(n\Delta t)\leq 1$, $\lambda^{\mathtt{CLIPER}}$ lies within $[0,1]$, yielding a valid convex combination between mean and current value. When $\rho(n\Delta t)<0$, the parameter may exceed unity, reflecting anti-correlated regimes where the mean contribution dominates. The complete derivation from $\mathtt{MSE}$ minimization is provided in B. Cyclostationary models provide a natural framework for processes exhibiting periodic statistical structure, allowing predictors to remain consistent with cyclic variations in mean and covariance.
@@ 2.2 Cyclostationary Extension of Forecasting Operators
Statistical Stationary Process Cyclostationary Process
$E[X(t)]$ $\mu(t)=\mu,\ \forall t$ $\mu(t)=\mu(t+T),\ \forall t$
Theorem 1 (Periodically Stationary or Cyclostationary Processes [17, 31]).
@@
In contrast to weakly stationary processes, cyclostationary signals preserve their statistical structure modulo the period $T$. Their moments evolve periodically, reflecting temporal symmetries common to many natural and engineered systems such as irradiance, wind speed, electrical load, or tidal dynamics. This property defines the statistical foundation required for analytical forecasting under periodic non stationarity.
@@
The forecasting workflow using $\mathtt{P}_{\mathtt{BLEND}}$ is summarized in Figure 1, which distinguishes the stationary, periodically stationary, and simplified analytical forms. The mathematical properties detailed in D ensure convexity, boundedness, and stability, core guarantees for applying $\mathtt{BLEND}$ operators to cyclostationary energy processes where periodicity and noise coexist.